A body starts slipping down on an incline plane and moves half meter in half second .How long will it take to next half metre?

Grade:12

2 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavi

motion for 1/2 meter  

so s= ut +1/2 a t2

 1/2 =0+1/2 * a *(1/2)2

 a=4 m/sec2

now motion for 1 meter

 s= ut +1/2 a t2

 1=0+ 1/2 * 4 *t2

t=1/ √2

so time taken for next 1/2 meter is  t'=1/ √2   -1/2  sec

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin

Vineeth Chennabhoina
7 years ago
First we find the acceleration of the body. For the first half second, the distance covered is half a meter. As we solve s= ut + 1/2at2, we get a=4m/s2. Intially the body didn’t have any velocity, but as we go through the first half second, the body gains a velocity v= u +at , as this is the velocity the body gains at the end of first half second and it has no initial velocity, substituting a=4 and t=0.5, we get v=2 m/s.
 
Finally, the time taken for the second half second will be
 
s = ut + ½at2. Here s is 0.5, u=2m/s, a=4m/s. We get a quadratic in t. Which would be
4t2 + 4t – 1 = 0 Solving by the finding roots of x, we get 1.656/8, which is 0.207 seconds, rounding it off to 0.21 seconds.